Molecular weight data for some polymer are tabulated. Determine:
a) The number-average molecular weight
b) The weight-average molecular weight
c) What is the number-average degree of polymerization for this polymer?
d) What is the weight-average degree of polymerization for this polymer?

Molecular weight data for some polymer are tabulatSide note: Please consider the monomer C3H6
Answer #1

Given:

              

Molecular Weight Range

(g/mol)

xi

wi

8,000 – 16,000

0.05

0.02

16,000 – 24,000

0.16

0.10

24,000 – 32,000

0.24

0.20

32,000 – 40,000

0.28

0.30

40,000 – 48,000

0.20

0.27

48,000 – 56,000

0.07

0.11

Solution:

           (a) From the tabulated data, let us find out Mn, the number-average molecular weight.

Molecular Weight Range

(g/mol)

Mean Mi

xi

xiMi

8,000 – 16,000

12,000

0.05

600

16,000 – 24,000

20,000

0.16

3,200

24,000 – 32,000

28,000

0.24

6,720

32,000 – 40,000

36,000

0.28

10,080

40,000 – 48,000

44,000

0.20

8,800

48,000 – 56,000

52,000

0.07

3,640

      Mn = \sumxiMi= 33,040 g/mol

             (b) From the tabulated data, let us find out Mw, the weight-average molecular weight.

Molecular Weight Range

(g/mol)

Mean Mi

wi

wiMi

8,000 – 16,000

12,000

0.02

240

16,000 – 24,000

20,000

0.10

2,000

24,000 – 32,000

28,000

0.20

5,600

32,000 – 40,000

36,000

0.30

10,800

40,000 – 48,000

44,000

0.27

11,880

48,000 – 56,000

52,000

0.11

5,720

      Mw = \sumxiMi= 36,240 g/mol

   We can define the average degree of polymerization (# of monomer units per polymer molecule) according to the number-average (nn) or weight-average (nn):

                             nn = Mn / m    or

                             nw = Mn / m

where m is the molecular weight of the monomer unit (in this case, C3H6).

(c) For number - average degree of polymerization:

                 Molecular weight of the monomer (m) = 3*12.011 + 6*1.00794

                                                                                 = 42.08 g/mol

                 Now, nn = Mn / m   

                 For which Mn = 33,040 g/mol

                 Now, calculating nn, we get,

                              

                           nn = 33,040 / 42.08 = 785

                           

(d) For weight - average degree of polymerization:

                 Molecular weight of the monomer (m) = 3*12.011 + 6*1.00794

                                                                                 = 42.08 g/mol

                 Now, nw = Mn / m   

                 For which Mn = 36,240 g/mol

                 Now, calculating nw, we get,

                              

                           nw = 36,240 / 42.08 = 861